A liquid at room temperature (27degree C) is poured to the brim of a glass flask whose volume is 1 liter. The flask + liquid are heated to 40degree C. How much liquid will spill over? The volume expansion coefficient of liquid is 75*10^(-5)/degreeC and that of the glass is 1.5*10^(-5)/degreeC
Answer #1

The expansion in volume of the liquid leads to change in volume

This change is ΔV = V1 ( γΔT ) where γ = coefficient of volume expansion of the liquid

ΔV = 1000( 75 * 10^-5 * ( 40-27 ) ) = 9.75 cm^3 [Since 1L = 1000 cm^3]

For glass:

ΔVglass = V1 ( γglassΔT)

= 1000( 1.5 * 10^-5(40- 27))

= 0.195 cm^3

Thus, overflow = 9.75 – 0.195

= 9.555 cm^3

Kindly upvote :)